OpenStudy (anonymous):

x^2-6x+9/12-4x * x^3-3x^2/x^6-9x^4 x^2-6x+9/12-4x * x^3-3x^2/x^6-9x^4 @Mathematics

OpenStudy (aroub):

x^2-6x+9 x^3 -3x^2 ---------- * ----------- 12-4x x^6-9x^4 = (x-3) (x-3) x^2 (x-3) --------- * ------------ 4(x-3) x^4 ( x+3)(x-3) = 1 x^2 (x-3) ---- * ---------- 4 x+3 = x^2(x-3) ---------- 4(x+3) \[x \neq -3 \] & \[x \neq 3\]

OpenStudy (anonymous):

thanks sooo much aroub :)

OpenStudy (aroub):

no problem :)